# REACT: Real-time Efficiency and Accuracy Compromise for Tradeoffs in Scene Graph Generation

### v2

@@ 4 Toward Real-Time SGG
Based on observations in Section 3, we propose a novel architecture for real-time SGG, which is depicted in Figure 2(c) and Figure 2(d). We apply the Decoupled Two-Stage architecture and replace the Faster-RCNN detector of PE-NET with a YOLO backbone, as shown in Figure 2(c). We improve the Prototype Learning of PE-NET by removing superfluous components, which results in Efficient Prototype Learning. We name the combination of DTS and Efficient Prototype Learning as the REACT model. We also propose Dynamic Candidate Selection (DCS) to further reduce the latency of the REACT model at inference time. In the following, we detail these components.
4.1 Decoupled Two-Stage SGG
@@
To maintain the object detector accuracy during relation prediction training, we propose to freeze the regression and the classification head of the object detector and perform NMS before the relation prediction stage. With this strategy, the object classifier in the relation prediction stage is also removed and class prediction from the object detector is used for the final prediction. The objective of the relation prediction stage becomes then to predict only the pairs and associate predicates, in contrast to also predicting the class labels of objects. The original loss used during relation prediction: $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{rel}}+\mathcal{L}_{\text{obj}}$ becomes: $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{rel}}$. This strategy significantly lowers the complexity and thus the computational load of the model. In addition, this approach helps to fairly evaluate the context learning of SGG models, independently of the object detector used. With our Decoupled Two-Stage architecture, any object detector can be used in place of Faster-RCNN, without impacting stage 2 of the architecture. For instance, one could consider replacing Faster-RCNN with a real-time object detector such as YOLO [11]. Recent versions of YOLO, such as YOLOV8 [15], YOLOV9 [39] or YOLOV10 [38] are very efficient for real-time object detection. To take advantage of YOLO versions, we leveraged the introduced Decoupled Two-Stage strategy and replaced the ResNeXt-101 backbone with YOLO’s CSPNet-based backbones [15, 39] (see Figure 2(c) (a)). P5-FPN of Faster-RCNN was also replaced by the YOLOV8/9/10 PAN-FPN Necks (see Figure 2(c) (b)). Finally, we removed the RPN and detection head of Faster-RCNN and replaced them with the YOLOV8/9/10 Detection Head (see Figure 2(c) (c)). Unlike Faster-RCNN, YOLO-based architectures are not designed for feature extraction because they do not implement any ROI Align module. To address this issue, we combined the feature maps extracted from the P3, P4, and P5 layers of YOLO and added one 3x3 Conv block to reduce the feature sizes before ROI Align (see Figure 2(c) (f)). The ROI Align module from Faster-RCNN has been modified to align features with corresponding bounding boxes at only three different scales (instead of five in the baseline implementation, see Figure 2(c) (f)). Thanks to our DTS strategy, switching from one YOLO version to another does not impact the SGG architecture.
4.2 Dynamic Candidate Selection
@@
Here, we introduce the Dynamic Candidate Selection (DCS) method to select an optimal number of proposal candidates at inference time. During training, the relation prediction stage takes as input a maximum amount of proposals $\theta$ ($\theta=100$). During evaluation, we uniformly sample the number of proposals $k$, starting from 0. Using the generated slope for each metric, we can compute the optimal Dynamic Candidate Selection threshold as follows:
$x_{\text{opt}}=\min\left\{x\mid\left|f^{\prime}(x)\right|<\epsilon\right\}$ (2)
During inference, $x_{\text{opt}}$ is applied to control the maximum number of proposals to use as input for the relation prediction stage, successfully lowering the complexity of the task.
4.3 Efficient Prototype Learning
@@
In PE-NET, node, and edge features are upsampled 8x and 4x respectively to a fixed size of 4096 by a feed-forward layer (see steps (g) and (h) in Figure 2(b)). We believe this is unnecessary and that the model will perform similarly without upsampling. We then removed this layer and kept the original feature sizes of 256 and 512, respectively. In the latter stage during prototype learning, we removed the associated downsampling layers. For the PE-NET model, the use of union features for the edge representation seems redundant with the visual features of the subject and object. As a result, we removed the use of union features and dedicated ROI Align step (see Figure 2(a) (d)). By doing so, the edge representation refinement step (see Figure 2(a) (g)) can also be removed and the spatial features fed directly to the context learning (see Figure 2(d) (i)). To further boost accuracy, we proposed to modify the fusion of subject $s$ and object $o$ features during Prototype Learning of the PE-NET model. We replaced the fusion operation: $ReLU(s+o)-(s-o)^{2}$ with a single linear layer and removed the subsequent gate operation.
Table 2: Full results of our experiments with the YOLOV8 and Faster-RCNN models on the test set of the PSG dataset. $\dagger$ represents evaluation with our DCS strategy. Bold and underlined represent the best and second best for each metric. $\downarrow$ and $\uparrow$ represent the improvements of DTS-YOLO against corresponding Two-Stage baselines with Faster-RCNN. See Figure 1 for a visual representation.
@@ B D Relation Head mR@20/50/100 R@20/50/100 F1@K mAP50 Latency Params
CSPDarkNet-53 YOLOV8m-DTS REACT (ours) $\dagger$  18.3 /  20.1 / 20.8 27.5 / 30.9 / 32.2 23.9 53.1 23.0 43.3M
REACT (ours) 18.3 / 20.0 / 20.9 27.6 / 30.9 / 32.3 23.9 53.1 32.5 43.3M
PE-NET $\dagger$ [46] 17.1 / 19.0 / 19.8 28.0 / 31.3 / 32.9 23.2 $\uparrow$9.1 53.1 $\uparrow$17.6 51.7 $\downarrow$338 187M
PE-NET [46] 17.1 / 19.0 / 19.9 28.0 / 31.3 / 33.0 23.2 $\uparrow$9.1 53.1 $\uparrow$17.6 103.0 $\downarrow$261 187M
GPS-NET $\dagger$ [26] 10.9 / 12.6 / 13.5 26.3 / 30.0 / 31.8 17.3 $\uparrow$5.7 53.1 $\uparrow$21.5 27.8 $\downarrow$286 45.3M
GPS-NET [26] 11.1 / 12.7 / 13.7 26.6 / 30.1 / 31.9 17.6 $\uparrow$5.9 53.1 $\uparrow$21.5 41.1 $\downarrow$272 45.3M
@@
Neural-Motifs $\dagger$ [44] 10.1 / 11.6 / 12.5 25.7 / 28.7 / 30.3 16.2 $\uparrow$3.8 53.1 $\uparrow$17.2 22.4 $\downarrow$296 52.9M
@@
Neural-Motifs [44] 10.1 / 11.6 / 12.5 25.7 / 28.7 / 30.4 16.2 $\uparrow$3.8 53.1 $\uparrow$17.2 35.8 $\downarrow$282 52.9M
VCTree $\dagger$ [34] 10.7 / 12.2 / 12.9 22.1 / 25.3 / 26.8 16.1 $\uparrow$3.7 53.1 $\uparrow$20.5 183.2 $\downarrow$336 268.0M
VCTree [34] 10.7 / 12.2 / 13.0 22.1 / 25.3 / 26.8 16.1 $\uparrow$3.8 53.1 $\uparrow$20.5 239.5 $\downarrow$280 268.0M
Transformer $\dagger$ [35] 11.3 / 12.4 / 12.8 24.1 / 27.0 / 28.3 16.6 $\uparrow$5.5 53.1 $\uparrow$18.5 21.1 $\downarrow$275 97.9M
Transformer [35] 11.3 / 12.4 / 13.2 24.1 / 27.1 / 28.9 16.9 $\uparrow$5.7 53.1 $\uparrow$18.5 35.1 $\downarrow$263 97.9M
ResNeXt-101 Faster-RCNN PE-NET [46] 10.7 / 11.7 / 12.0 16.8 / 18.7 / 19.5 14.1 35.5 390.4 426.5M
GPS-NET [26] 8.2 / 8.8 / 9.2 15.7 / 18.1 / 19.2 11.7 31.6 313.9 391.6M
Neural-Motifs [44] 8.9 / 9.6 / 9.8 16.9 / 18.7 / 19.5 12.5 35.9 318.6 369.6M
VCTree [34] 7.9 / 9.1 / 9.6 18.0 / 20.7 / 22.3 12.4 32.6 519.5 365.9M
Transformer [35] 7.6 / 8.2 / 8.5 16.2 / 17.9 / 18.7 11.1 34.6 296.2 328.3M
DETR EGTR [10] 12.0 / 14.5 / 16.6 24.9 / 30.3 / 33.8 19.4 33.6 78.3 42.5M
RelTR [4] 7.9 / 10.0 / 11.7 12.9 / 17.3 / 20.2 12.4 25.7 59.6 63.7M
5 Experiments
Table 3: Full results of our experiments on the test set of IndoorVG. mR@K and R@K represent the average for $k=[20,50,100]$. $\dagger$ represents evaluation with our DCS strategy.
@@ B Relation Head mR@K R@K F1@K mAP50
Transformer 9.9 16.8 12.5 24.3
OS EGTR [10] 7.1 10.6 8.5 14.7
RelTR [4] 7.7 11.4 9.2 14.1
Datasets. In SGG, the most widely used dataset is Visual Genome [17]. The commonly used split of the data is called VG150 and contains annotations from the top 150 object classes and top 50 predicate classes. However, VG150 suffers from severe biases [42, 29, 8] and inherent issues such as the presence of ambiguous classes (e.g. the classes ”people”, ”men” etc) [42, 24]. Recently, two new datasets have been proposed with higher-quality annotations: the PSG [42] and IndoorVG [28] datasets. Both datasets have been carefully annotated to remove the presence of ambiguous classes or wrong annotations, which makes them a better choice for benchmarking our approach.
Metrics. Following previous work [41, 34], we used the Recall@K (R@K) and meanRecall@K (mR@K) metrics to measure the performance of models. For both metrics, we evaluate the recall on the top $K$ ($k=[20,50,100]$) relations predicted, ranked by confidence. The Recall@K evaluates the overall performance of a model on the selected dataset whereas meanRecall@K evaluates the performance on the average of all predicate classes, which is more significant for long-tail learning such as in the task of SGG. Both metrics are averaged in the F1@K metric as follows: $F1@K=\frac{2\times R@K\times mR@K}{R@K+mR@K}$ which efficiently represents the trade-off for the model performance between head and tail predictions. Object Detection (OD) accuracy is measured using standard mAP@50 (mAP50).
Implementation details. The latest YOLO versions (V8, V9, and V10) can be used interchangeably in our approach. For simplicity here, we will present experiments for the YOLOV8 medium version [15] (YOLOV8m). YOLOV8m has been trained for object detection for 100 epochs on all datasets with a batch size of 16 and image size 640x640, following the official implementation [^2] https://github.com/ultralytics/ultralytics. All relation prediction models have been trained for 20 epochs, batch size 8 with the SGD optimizer, and an initial learning rate of $1e^{-3}$. For Dynamic Candidate Selection (DCS), we choose $\epsilon=1e^{-5}$. Inference without DCS is performed with a fixed number of 100 proposals for all models. In contrast to YOLO, models trained with Faster-RCNN use 600x1000 image sizes. We fully re-trained seven different relation prediction models on the PSG and IndoorVG datasets: Neural-Motifs [44], VCTree [34], Transformer [35], GPS-Net [26], PE-NET [46], RelTR [4] and EGTR [9] following provided codebases. All models are implemented without any re-weighting [37] or de-biasing strategy [35]. Latency is benchmarked with a batch size of 1333Hardware used: 11th Gen IntelTM CoreTM i9-11950H @ 2.60GHz x 16, NVIDIA GeForce RTX 3080 Laptop GPU 16GB VRAM, 2 x 16GB 3200 MHz RAM..
[…]
Table 4: REACT compared to previous work on VG150 [41, 17] (reported results). mR@K and R@K are the average for $k=[50,100]$ as some approaches are not reporting $k=20$.

### v3

@@ 4 Toward Real-Time SGG
This work proposes a novel architecture for real-time SGG, as shown in Figure˜2(c) and Figure˜2(d). This architecture applies the Decoupled Two-Stage method, replacing the Faster-RCNN detector of PE-NET with a YOLO backbone (cf. Figure˜2(c)). The Prototype Learning of PE-NET is improved by removing superfluous components, which results in Efficient Prototype Learning. This combination of DTS and Efficient Prototype Learning constitutes the REACT model. We also propose Dynamic Candidate Selection (DCS) to further reduce the latency of the proposed model at inference time, as described below.
@@
Decoupled Two-Stage SGG. To maintain the object detector accuracy during relation prediction training, we propose to freeze the regression and the classification head of the object detector and perform NMS before the relation prediction stage. With this strategy, the object classifier in the relation prediction stage is also removed and class prediction from the object detector is used for the final prediction. Thus, the relation prediction stage outputs only the pairs and associate predicates, in contrast to also predicting the class labels of objects. The loss used during relation prediction: $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{rel}}+\mathcal{L}_{\text{obj}}$, then becomes: $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{rel}}$, lowering the model complexity and, consequently, its computational load. This approach helps to evaluate the context learning of SGG models independently of the object detector used. With this architecture, any object detector can be used in place of Faster-RCNN, without impacting stage 2 of the architecture. For instance, one could consider replacing Faster-RCNN with a real-time object detector such as YOLO [Jin et al.(2023a)Jin, Guo, Meng, Zhu, Xi, Wang, Mu, and Song]. To take advantage of YOLO versions, we leveraged the introduced Decoupled Two-Stage strategy and replaced the ResNeXt-101 backbone with YOLO’s CSPNet-based backbones [Jocher et al.(2023)Jocher, Chaurasia, and Qiu, Wang et al.(2024b)Wang, Yeh, and Liao] (Figure˜2(c) (a)). P5-FPN of Faster-RCNN was also replaced by the YOLOV8/9/10 PAN-FPN Necks (Figure˜2(c) (b)). Finally, we removed the RPN and detection head of Faster-RCNN and replaced them with the YOLOV8/9/10 Detection Head (Figure˜2(c) (c)). Unlike Faster-RCNN, YOLO-based architectures are not designed for feature extraction because they do not implement any ROI Align module. To address this issue, we combined the feature maps extracted from the P3, P4, and P5 layers of YOLO and added one 3x3 Conv block to reduce the feature sizes before ROI Align (Figure˜2(c) (f)). The ROI Align module from Faster-RCNN has been modified to align features with corresponding bounding boxes at only three different scales (instead of five in the baseline implementation, Figure˜2(c) (f)).
@@
Dynamic Candidate Selection. Here, we introduce the DCS method to select an optimal number of proposal candidates at inference time. During training, the relation prediction stage takes as input a maximum amount of proposals $\theta$ ($\theta=100$). During evaluation, we uniformly sample the number of proposals $k$, starting from 0. Using the generated slope for each metric, the optimal DCS threshold is computed as : $x_{\text{opt}}=\min\left\{x\mid\left|f^{\prime}(x)\right|<\epsilon\right\}.$ During inference, $x_{\text{opt}}$ is applied to control the maximum number of proposals to use as input for the relation prediction stage, successfully lowering the complexity of the task.
@@
Efficient Prototype Learning. In PE-NET, node, and edge features are upsampled 8x and 4x respectively to a fixed size of 4096 by a feed-forward layer (steps (g) and (h) in Figure˜2(b)). We believe this is unnecessary and that the model will perform similarly without upsampling. We then removed this layer and kept the original feature sizes of 256 and 512, respectively. In the latter stage during prototype learning, we removed the associated downsampling layers. For the PE-NET model, the use of union features for the edge representation seems redundant with the visual features of the subject and object. As a result, we removed the use of union features and dedicated ROI Align step (Figure˜2(a) (d)). By doing so, the edge representation refinement step (Figure˜2(a) (g)) can also be removed and the spatial features fed directly to the context learning (Figure˜2(d) (i)). To further boost accuracy, we proposed to modify the fusion of subject $s$ and object $o$ features during Prototype Learning of the PE-NET model. We replaced the fusion operation: $ReLU(s+o)-(s-o)^{2}$ with a single linear layer and removed the subsequent gate operation.
5 Experiments
@@ B D Relation Head mR@20/50/100 R@20/50/100 F1@K mAP50 Latency Params
CSPDarkNet-53
YOLOV8m-DTS
REACT (ours) $\dagger$ 18.3 / 20.0 / 20.8 27.5 / 30.9 / 32.2 23.9 53.1 23.0 43.3M
REACT (ours) 18.3 / 20.1 / 20.9 27.6 / 30.9 / 32.3 23.9 53.1 32.5 43.3M
PE-NET $\dagger$ [Zheng et al.(2023)Zheng, Lyu, Gao, Dai, and Song] 17.1 / 19.0 / 19.8 28.0 / 31.3 / 32.9 23.2 $\uparrow$9.1 53.1 $\uparrow$17.6 51.7 $\downarrow$338 187M
PE-NET [Zheng et al.(2023)Zheng, Lyu, Gao, Dai, and Song] 17.1 / 19.0 / 19.9 28.0 / 31.3 / 33.0 23.2 $\uparrow$9.1 53.1 $\uparrow$17.6 103.0 $\downarrow$261 187M
GPS-NET $\dagger$ [Lin et al.(2020)Lin, Ding, Zeng, and Tao] 10.9 / 12.6 / 13.5 26.3 / 30.0 / 31.8 17.3 $\uparrow$5.7 53.1 $\uparrow$21.5 27.8 $\downarrow$286 45.3M
GPS-NET [Lin et al.(2020)Lin, Ding, Zeng, and Tao] 11.1 / 12.7 / 13.7 26.6 / 30.1 / 31.9 17.6 $\uparrow$5.9 53.1 $\uparrow$21.5 41.1 $\downarrow$272 45.3M
@@
Neural-Motifs $\dagger$ [Zellers et al.(2018)Zellers, Yatskar, Thomson, and Choi] 10.1 / 11.6 / 12.5 25.7 / 28.7 / 30.3 16.2 $\uparrow$3.8 53.1 $\uparrow$17.2 22.4 $\downarrow$296 52.9M
@@
Neural-Motifs [Zellers et al.(2018)Zellers, Yatskar, Thomson, and Choi] 10.1 / 11.6 / 12.5 25.7 / 28.7 / 30.4 16.2 $\uparrow$3.8 53.1 $\uparrow$17.2 35.8 $\downarrow$282 52.9M
VCTree $\dagger$ [Tang et al.(2019)Tang, Zhang, Wu, Luo, and Liu] 10.7 / 12.2 / 12.9 22.1 / 25.3 / 26.8 16.1 $\uparrow$3.7 53.1 $\uparrow$20.5 183.2 $\downarrow$336 268.0M
VCTree [Tang et al.(2019)Tang, Zhang, Wu, Luo, and Liu] 10.7 / 12.2 / 13.0 22.1 / 25.3 / 26.8 16.1 $\uparrow$3.8 53.1 $\uparrow$20.5 239.5 $\downarrow$280 268.0M
Transformer $\dagger$ [Tang et al.(2020)Tang, Niu, Huang, Shi, and Zhang] 11.3 / 12.4 / 12.8 24.1 / 27.0 / 28.3 16.6 $\uparrow$5.5 53.1 $\uparrow$18.5 21.1 $\downarrow$275 97.9M
Transformer [Tang et al.(2020)Tang, Niu, Huang, Shi, and Zhang] 11.3 / 12.4 / 13.2 24.1 / 27.1 / 28.9 16.9 $\uparrow$5.7 53.1 $\uparrow$18.5 35.1 $\downarrow$263 97.9M
ResNeXt-101
Faster-RCNN
PE-NET [Zheng et al.(2023)Zheng, Lyu, Gao, Dai, and Song] 10.7 / 11.7 / 12.0 16.8 / 18.7 / 19.5 14.1 35.5 390.4 426.5M
GPS-NET [Lin et al.(2020)Lin, Ding, Zeng, and Tao] 8.2 / 8.8 / 9.2 15.7 / 18.1 / 19.2 11.7 31.6 313.9 391.6M
Neural-Motifs [Zellers et al.(2018)Zellers, Yatskar, Thomson, and Choi] 8.9 / 9.6 / 9.8 16.9 / 18.7 / 19.5 12.5 35.9 318.6 369.6M
VCTree [Tang et al.(2019)Tang, Zhang, Wu, Luo, and Liu] 7.9 / 9.1 / 9.6 18.0 / 20.7 / 22.3 12.4 32.6 519.5 365.9M
Transformer [Tang et al.(2020)Tang, Niu, Huang, Shi, and Zhang] 7.6 / 8.2 / 8.5 16.2 / 17.9 / 18.7 11.1 34.6 296.2 328.3M
DETR
EGTR [Im et al.(2024b)Im, Nam, Park, Lee, and Park] 12.0 / 14.5 / 16.6 24.9 / 30.3 / 33.8 19.4 33.6 78.3 42.5M
RelTR [Cong et al.(2022b)Cong, Yang, and Rosenhahn] 7.9 / 10.0 / 11.7 12.9 / 17.3 / 20.2 12.4 25.7 59.6 63.7M
Table 2: Results on the PSG dataset. $\dagger$ represents evaluation with our DCS strategy. Bold and underlined represent the best and second best for each metric. $\downarrow$ and $\uparrow$ represent the improvements of DTS-YOLO against the corresponding baselines.
Datasets. In SGG, the most widely used dataset is Visual Genome [Krishna et al.(2017)Krishna, Zhu, Groth, Johnson, Hata, Kravitz, Chen, Kalantidis, Li, Shamma, Bernstein, and Fei-Fei]. The commonly used split of the data is called VG150 and contains annotations from the top 150 object classes and top 50 predicate classes. However, VG150 suffers from severe biases [Yang et al.(2022a)Yang, Ang, Guo, Zhou, Zhang, and Liu, Neau et al.(2023b)Neau, Santos, Bosser, and Buche, Hudson and Manning(2019)] and inherent issues such as the presence of ambiguous classes (e.g. the classes "people", "men" etc) [Yang et al.(2022a)Yang, Ang, Guo, Zhou, Zhang, and Liu, Liang et al.(2019)Liang, Bai, Zhang, Qian, Zhu, and Mei]. Recently, two new datasets have been proposed with higher-quality annotations: the PSG [Yang et al.(2022a)Yang, Ang, Guo, Zhou, Zhang, and Liu] and IndoorVG [Neau et al.(2023a)Neau, Santos, Bosser, and Buche] datasets. Both datasets have been carefully annotated to remove the presence of ambiguous classes or wrong annotations, which makes them a better choice for benchmarking our approach.
Metrics. Following previous work [Xu et al.(2017)Xu, Zhu, Choy, and Fei-Fei, Tang et al.(2019)Tang, Zhang, Wu, Luo, and Liu], we used the Recall@K (R@K) and meanRecall@K (mR@K) metrics to measure the performance of models. For both metrics, we evaluate the recall on the top $K$ ($k=[20,50,100]$) relations predicted, ranked by confidence. The Recall@K evaluates the overall performance of a model on the selected dataset whereas meanRecall@K evaluates the performance on the average of all predicate classes, which is more significant for long-tail learning such as in the task of SGG. Both metrics are averaged in the F1@K metric as follows: $F1@K=\frac{2\times R@K\times mR@K}{R@K+mR@K}$ which efficiently represents the trade-off for the model performance between head and tail predictions. Object Detection (OD) accuracy is measured using standard mAP@50 (mAP50).
Implementation details. The latest YOLO versions (V8, V9, and V10) can be used interchangeably in our approach. For simplicity here, we will present experiments for the YOLOV8 medium version [Jocher et al.(2023)Jocher, Chaurasia, and Qiu] (YOLOV8m). YOLOV8m has been trained for object detection for 100 epochs on all datasets with a batch size of 16 and image size 640x640. Relation prediction models have been trained for 20 epochs, batch size 8 with the SGD optimizer, and an initial learning rate of $1e^{-3}$. For Dynamic Candidate Selection (DCS), we choose $\epsilon=1e^{-5}$. Inference without DCS is performed with 100 proposals for all models. In contrast to YOLO, models trained with Faster-RCNN use 600x1000 image sizes. We fully re-trained seven different relation prediction models on the PSG and IndoorVG datasets: Neural-Motifs [Zellers et al.(2018)Zellers, Yatskar, Thomson, and Choi], VCTree [Tang et al.(2019)Tang, Zhang, Wu, Luo, and Liu], Transformer [Tang et al.(2020)Tang, Niu, Huang, Shi, and Zhang], GPS-Net [Lin et al.(2020)Lin, Ding, Zeng, and Tao], PE-NET [Zheng et al.(2023)Zheng, Lyu, Gao, Dai, and Song], RelTR [Cong et al.(2022b)Cong, Yang, and Rosenhahn], and EGTR [Im et al.(2024a)Im, Nam, Park, Lee, and Park] following the provided codebases. All models are implemented without any re-weighting [Teng and Wang(2022b)] or de-biasing strategy [Tang et al.(2020)Tang, Niu, Huang, Shi, and Zhang]. Latency is benchmarked with a batch size of 1111Hardware used: 11th Gen IntelTM CoreTM i9-11950H @ 2.60GHz x 16, NVIDIA GeForce RTX 3080 Laptop GPU 16GB VRAM, 2 x 16GB 3200 MHz RAM..
Results. In Table˜2 and Table˜4 we display the performance of our DTS architecture in OD accuracy compared to previous work. The F1@K results from Table˜2 are represented as a graph in Figure˜1. In the text, we refer to improvements as mean absolute percentage.
@@ B Relation Head mR@K R@K F1@K mAP50
Transformer 9.9 16.8 12.5 24.2
OS
EGTR [Im et al.(2024b)Im, Nam, Park, Lee, and Park] 7.1 10.6 8.5 14.7
RelTR [Cong et al.(2022b)Cong, Yang, and Rosenhahn] 7.7 11.4 9.2 14.1
Table 3: Results on IndoorVG. mR@K and R@K are average for $k=[20,50,100]$.